OpenStudy (anonymous):

Eliminate the parameter x = t - 3, y = t2 + 5

OpenStudy (jango_in_dtown):

x=t-2 imply t=x+2 and hence y=t^2+5=(x+2)^2+5 imply x^2+4x+9-y=0

OpenStudy (anonymous):

Thank you @jango_IN_DTOWN but the answers are a.y = x2 + 6x + 14 b.y = x2 - 14 c. y = x2 - 6x - 14 d.y = x2 + 14

OpenStudy (jango_in_dtown):

oh sorry it is x=t-3 so t=x+3 and hence y=t^2+5=(x+3)^2+5=x^2+6x+14

OpenStudy (jango_in_dtown):

a. is correct..:)

OpenStudy (anonymous):

Thank you @jango_IN_DTOWN